OpenStudy (anonymous):

Solve the system by the elimination method. -2x + y + 6 = 0 2x + y - 8 = 0 When you eliminate x, what is the resulting equation? 2y = 2 2y = -2 y = -2

hartnn (hartnn):

Add those 2 equations, what do u get?

OpenStudy (anonymous):

y+-2

hartnn (hartnn):

+6-8 = -2 so 2y -2 =0 or 2y =2 :)

OpenStudy (anonymous):

Thank you! can you help me with a few more?

hartnn (hartnn):

sure :)

OpenStudy (anonymous):

Solve the system by the elimination method. Check your work. 3a + 5b - 7 = 0 a - 2b - 4 = 0

hartnn (hartnn):

so what have you tried? did you multiply second equation by 3 ?

OpenStudy (anonymous):

I have no idea how to even begin this lol. I just got switched to online and have no one to help me lol

hartnn (hartnn):

multiply second equation by 3 what do u get?

OpenStudy (anonymous):

3a-6b-12=0

hartnn (hartnn):

yep, now subtract that from the first equation! and you'll see 'a' getting eliminated :)

OpenStudy (anonymous):

I'm confused. These are the multiple choice answers. {(96/11, -5/11)} {(34/11, -5/11)} {(32/33, 5/11)}